OpenStudy (anonymous):

Solve for y. 4x+8y=17

OpenStudy (anonymous):

What's x? If we don't know what x is we can't solve for y

OpenStudy (anonymous):

it doesn't say what x is

OpenStudy (anonymous):

you have to move all of numbers to right side ,except y

OpenStudy (anonymous):

8y=17-4x y=(17-4x)/8

OpenStudy (anonymous):

8y = 17-4x y= 17/8 -1/2x

OpenStudy (anonymous):

oh ok thank u